OpenStudy (anonymous):

Solve each system using substitution. y = 4x -8 y =2x+10 Help please

OpenStudy (anonymous):

Nooooooo JESS PLIS

OpenStudy (anonymous):

</3

OpenStudy (sleepyjess):

You have to substitute one of the equations in for y So 4x - 8 = 2x + 10 Get the x on one side by subtracting 2x from each side 4x - 2x - 8 = 2x - 2x + 10 2x - 8 = 10

OpenStudy (sleepyjess):

Do you know what to do from here?

OpenStudy (sleepyjess):

@Godzilla123123

OpenStudy (anonymous):

Not really

OpenStudy (sleepyjess):

2x - 8 + 8 = 10 + 8

OpenStudy (sleepyjess):

2x = 18

OpenStudy (sleepyjess):

Then divide both sides by 2 2x/2 = 18/2 x = 9

OpenStudy (the_fizicx99):

y =2x+10 -2x +(4x -8) = 10 -2x + 4x - 8 = 10 2x = 18 x = 9 -2x + y = 10 -2(9) + y = 10 -18 + y = 10 +18 +18 y = 28 You can change one of them back if you want :P

OpenStudy (sleepyjess):

Then plug in the 9 for your x to get your y

OpenStudy (sleepyjess):

y = 4(9) - 8 y = 36 - 8 y = 28

OpenStudy (the_fizicx99):

Nerd off, you and me =.=

OpenStudy (sleepyjess):

to double check, do your 2nd equation also y = 2(9) + 10 y = 18 + 10 y = 28

OpenStudy (sleepyjess):

Do you understand all that @Godzilla123123 ?

OpenStudy (sleepyjess):

@tHe_FiZiCx99 ... what does that mean?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Yeah I do

OpenStudy (sleepyjess):

:-)

OpenStudy (anonymous):

thanks! <3

OpenStudy (the_fizicx99):

You and me, nerd vs nerd LOL xD

OpenStudy (sleepyjess):

LOL

OpenStudy (anonymous):

I'll give another problem :3

OpenStudy (sleepyjess):

Ok... @Godzilla123123 ... which explanation was better???

OpenStudy (anonymous):

Oh

OpenStudy (the_fizicx99):

*cough* mine *cough*

OpenStudy (sleepyjess):

NOT...hmmmmm

OpenStudy (the_fizicx99):

=.=

OpenStudy (sleepyjess):

I showed what to do with each side step by step

OpenStudy (sleepyjess):

:-)

OpenStudy (anonymous):

You win fizic

OpenStudy (anonymous):

she left

OpenStudy (the_fizicx99):

Well me too >.> | <.< | >.>

OpenStudy (sleepyjess):

What?!!

OpenStudy (sleepyjess):

I never left... just had to wait to finish answering

OpenStudy (the_fizicx99):

^^^^

OpenStudy (anonymous):

Oh

OpenStudy (sleepyjess):

And I posted what x was before @tHe_FiZiCx99 ... and I sent you a message explaining :-)

OpenStudy (anonymous):

Ik I would've picked you but I thought you left

OpenStudy (sleepyjess):

Besides, I'm prettier :-)

OpenStudy (anonymous):

Lol

OpenStudy (the_fizicx99):

Weeeeeeelll You've never seen this stud behind a computer all day ^~^ LMAO